OpenStudy (anonymous):

find 2 consecutive odd integers the sum of whose squares is 130

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Let x = the first odd integer Then x+2= second odd integer x^2=square of first odd integer (x+2)^2= square of second odd integer x^2+(x+2)^2=180

OpenStudy (mertsj):

*130 not 180

OpenStudy (primeralph):